OpenStudy (anonymous):

what number is 32 Percent of 75? Which of the following equations could be used to solve the problem? A. n= 0.32(75) B. 0.32n=75 C.n=75 divided by 0.32

OpenStudy (anonymous):

Do you know proportion method?

OpenStudy (anonymous):

part of the number/the number = percent over 100

OpenStudy (anonymous):

Now solve !

OpenStudy (anonymous):

@Tama34

OpenStudy (anonymous):

Okey! Thank you! <3

OpenStudy (anonymous):

again it is 35/x = 75/100

OpenStudy (anonymous):

sorry 32 not 35

OpenStudy (anonymous):

32/x = 75/100

OpenStudy (anonymous):

\[\frac{ 32 }{ x } = \frac{ 75 }{ 100 }\]

OpenStudy (anonymous):

Thank you for the help! :D It's really appreciated!

OpenStudy (anonymous):

Thanks! Kindly choose a best response and close the question!

OpenStudy (butterflydreamer):